OpenStudy (anonymous):

Fuel cells have been developed that can generate a large amount of energy. For example, a hydrogen fuel cell works by combining hydrogen and oxygen gas to produce water and electrical energy. If a fuel cell can generate 10.0 kilowatts of power and the current is 15.8 amps, what is the voltage of the electricity? A. 0.63 volts B. 158volts C. 633 volts D. 158,000 volts E. 5.8 volts

OpenStudy (souvik):

power P=V*I where V=voltage and I=current

OpenStudy (souvik):

I=P/V

OpenStudy (anonymous):

so its d

OpenStudy (anonymous):

is it B or D?

OpenStudy (souvik):

V=10,000/15.8

OpenStudy (anonymous):

c

OpenStudy (souvik):

right..

OpenStudy (anonymous):

lmao like I sadi Im in a class right now I failed this course

OpenStudy (anonymous):

P=VI....V=P/I....puting values...power in watts 10*1000=10,000. V=10,000/15.8=632.9...C is correct

OpenStudy (anonymous):

Ty to both of you

OpenStudy (anonymous):

:)